OpenStudy (anonymous):

Prove that the given equation is an identity. sin 3θ = 3 sin θ − 4 sin3 θ Working from the left-hand side, rewrite 3θ as 2θ + θ, and then use the addition formula for sine

OpenStudy (anonymous):

\[\sin(\alpha + \beta) =\sin\alpha\cos\beta + \sin\beta\cos\alpha\]